What is the value of k if one exact x-intercept of y=x^2+kx+1 is 3+2sqrt2 ?
Answer by stanbon(75887)

What is the value of k if one exact x-intercept of y=x^2+kx+1 is 3+2sqrt2 ?
-----------
If 3+2sqrt2 is a root f(3+2sqrt2)=0
----------
Therefore: (3+2sqrt2)^2 + k(3+2sqrt2)+1 = 0
k(3+2sqrt2)= -(3+2sqrt2)^2-1
k = -(3+2sqrt2) -1/(3+2sqrt2)
k = -(3+2sqrt2) - (3-2sqrt2)/(9-8)
k = -3-2sqrt2-3+2sqrt2
k = -6
==========
Cheers,
Stan H.